OpenStudy (anonymous):

Going back to the dog whistle in question 1, what is the minimum riding speed needed to be able to hear the whistle? Remember, you can assume the following things: The whistle you use to call your hunting dog has a frequency of 21.0 kHz, but your dog is ignoring it. You suspect the whistle may not be working, but you can't hear sounds above 20.0 kHz. The speed of sound is 330m/s at the current air temperature. I don't think "Question 1" is really needed in order to answer this question. I'm just really confused as to how I would approach this problem. Any help would be great. Thanks!

OpenStudy (anonymous):

@abb0t @ganeshie8 Would you be able to help me with this?

OpenStudy (rajat97):

i think we need that question1 to know the situation

OpenStudy (rajat97):

Or else, i'll give it a random try. I assume that we are going away from the whistle. so the frequency that we hear is of 20kHz(max. we can hear). This would require the minimum riding speed. we can use the concept of Doppler effect. it states f'=fo(vs +/- v1)/(vs +/- v2) and here, f' is the frequency heard, fo is the original frequency, vs is the speed of sound, v1 is the speed of observer and v2 is the speed of the source of sound. we use the numerator vs+v1 if the observer is moving towards the source of sound and vice-versa. We use the denominator vs+v2 when the source is moving away from the observer and vice-versa. in this question, f'=20kHz, fo=21kHz, vs=330m/s, v1=?(we are the observer), and v2=0(as the source is not moving) we use the numerator vs-v1 as we assumed that we are moving away from the source of the sound i.e. the whistle. So, by putting everything in the equation, i get v1=15.715m/s so the riding speed should be 15.715m/s i hope this helps you

OpenStudy (rajat97):

Thanks for the medal @IrishBoy123!